# Scalable Private Search with Wally

### v5

@@ 4.2 SHE details in keyword search
    First expand $\hat{q}$ to two $\sqrt{N/B}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. The server uses the oblivious expansion algorithm given in [4] to expand. This algorithm requires $2\sqrt{N/B}$ homomorphic substitutions.
@@
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/B$ plaintext-ciphertext multiplications and $N/B-\sqrt{N/B}$ ciphertext additions.
@@
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/B}$ ciphertext-ciphertext multiplications and $\sqrt{N/B}-1$ ciphertext additions. to compute.
@@ 6 Security
6.1 DP Security in the Central Model
@@
6.2 Generating view of the server
@@
To complete the proof, note that in Line 3, the simulator generates $\mathsf{list}$ in the same way as the curator in the central model. After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
7 Implementation and Performance Evaluation
In this section, we demonstrate Wally’s concrete performance via a set of experiments. We also showcase Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
7.1 Implementation
We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [6]. Following the library, we use BFV [17] to implement the SHE parts of Wally. However, our protocol could easily be implemented using other SHE schemes such as BGV [10].
BFV parameters
The choice of BFV parameters provides a trade-off between computation and communication in Wally. At a high level, larger parameters provide better computation but result in high request size. We tried multiple sets of parameters and picked the one that provided the best performance. We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits. These parameters allow a precision of six bits. To increase the precision, we use plaintext CRT optimization described in Section A. This optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query). For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$. This gives us a precision of 13 bits.
Differential privacy parameters
Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries. For a fixed failure probability $\delta$, smaller $\epsilon$ provides stronger privacy but worse performance because each client is required to make more fake queries. We set $\delta=2^{-20}$, which means DP will fail only once a million and $\epsilon=1$, providing a decent privacy guarantee.
7.2 Performance Evaluation
@@
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Query Size (MB) 0.67 0.84 16.2
Response Size
(MB) 0.16 0.18 3.2
Time to process
queries 117 mins 139 mins 2.8 years
Precision (bits) 6 13 4
Privacy DP DP SHE
Table 2: Performance comparison of Wally and Tiptoe for database of 35 millions entries. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8 million real queries, or 24 million in total. All Wally responses are computed as the encrypted scores plus $44$ KB of metadata.
Table 2 describes our main experiments compared to Tiptoe. For an end-to-end system, we estimate the network time on a 30Mbps channel. This comes out to 74ms for transmitting a query and response over the network for the small precision setting (first column in Tables 2 and 3).
The server database consists of 35 million entries divided into around 8500 clusters. We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$. We assume each client makes $\Delta=2$ queries per epoch for these experiments. Moreover, there are four million honest clients present in an epoch. This results in eight million real queries in an epoch.
For Wally, we calculated the overhead by using Theorem 2. The fake query overhead is derived by the expected number of fake queries for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
where $K$ is the total number of clusters and $U$ is the total number of clients. For our selected parameters, for each real query, the client is required to make just under two fake queries. Therefore, the total number of queries within an epoch is 24 million. In Table 2, we included fake query overhead in query request, response and total computation plus estimated network time in Wally. The epoch size in Wally is the total time it takes to process all the queries (real and fake).
For Tiptoe, we assume there are eight million queries the server processes one by one. Tiptoe requires running experiments on real compute nodes; due to resource limitations, we carefully estimate their results using Table 7 in [25].
Before discussing Wally’s performance, we emphasize that Wally requires the following two extra assumptions not required in Tiptoe.
  • Wally requires sufficiently many honest clients during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
In other words, Wally trades off privacy guarantees for scalability.
Wally has significantly better performance than Tiptoe across all the metrics. Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe. This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters. Similarly, the response size is 3.1-3.6 smaller. Wally’s efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe. For example, to process eight million queries, the server with 320 cores requires around 117 minutes in Wally. In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
Scaling DP with $\epsilon$.
Figure 4 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients. This is derived by dividing Equation 1 by the number of clusters $K$. Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
Figure 4: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section A. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 19 35 110
Table 3: Here we list the query and response size per each query (real and fake). Each response size includes $44$kB of metadata. Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
Scaling with the number of clients.
Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system. We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients. As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads. Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
Number of Clients .5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$ $0.65$ $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$ $0.61$
Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
7.3 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [6]. Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4]. Each response size was 27KB. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
Database entries Entry Size (Bytes) Query size (KB) Time (ms)
100k 10 368 166
100k 100 539 465
1 million 10 539 565
1 million 100 624 1700
Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section A.
8 Related Work
Anonymization networks alone do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness. Tiptoe [25], the most recent and performant, has been discussed above in-detail. Other less efficient solutions include Coeus [1] as well as Hers [16], with the latter providing privacy for the server’s database as well. Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[7, 31], and including a client public key in each query [19].

### v6

@@ IV-B Metadata Fetch
    First expand $\hat{q}$ to two $\sqrt{N^{\prime}}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. The server uses the oblivious expansion algorithm given in [33] to expand. This algorithm requires $\sqrt{N^{\prime}}$ homomorphic substitutions.
@@
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N$ plaintext-ciphertext multiplications and ciphertext additions.
@@
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N^{\prime}}$ ciphertext-ciphertext multiplications and additions.
@@ VI Security
VI-A DP Security in the Central Model
@@
VI-B Generating view of the server
@@
To complete the proof, observe that in Line 3 of Hybrid 2, the simulator generates $\mathsf{list}$ in a manner similar to the curator in the central model. However, the key difference lies in the scheduling of queries. The simulator sends queries from $\mathsf{list}$ at random intervals, while the curator in the central model sends $\mathsf{list}$ all at once to the server. Despite the adversary potentially recognizing its interaction with the simulator, it gains no additional advantage because the random schedule remains independent of secret inputs from the user.
VII Implementation
We have implemented Wally in Swift and Python languages. We next discuss implementation details for each component of Wally
VII-A Open Source SHE library
Components of Wally outlined in Section IV are instantiated using the Brakerski-Fan-Vercauteren (BFV) encryption scheme [17]. We have released the SHE components of Wally as an open-source Swift library named Apple Swift Homomorphic Encryption Library available at https://github.com/apple/swift-homomorphic-encryption. It includes the BFV scheme, a secure dot-product for search, and Keyword PIR for metadata fetch. We have implemented many optimizations within the library, which significantly improve the overall performance and can be of independent interest. The library API is simple and flexible, and can be used for other applications. We also provide an easy to follow tutorial and examples to use the API.
We discuss a few key optimizations and implementation decisions:
  • Residue Number System (RNS) variant of BFV. In BFV scheme, we manipulate elements modulo ciphertext modulus $Q$, which is an integer with hundreds of bits. Implementing this scheme requires slow multi-precision modulo arithmetic. To address this, we instead implemented the RNS variant of BFV [35]. In this variant, $Q$ is chosen as $Q=q_{1},\ldots,q_{l}$, and each $q_{i}$ is a single-precision integer of typically 28 or 55 bits. By using the RNS, any integer $x\in Z_{Q}$ can be represented as $\{x_{1}=x\mod q_{1},\cdots,x_{l}=x\mod q_{l}\}$ and the operation on $x\mod Q$ can be performed by performing the same operation on each RNS component $x_{i}$ modulus $q_{i}$. At the end we can use CRT to extract original ciphertext in $Z_{Q}$. We also implemented the Hybrid RNS variant of key-switching [36] and fast base conversion [37] methods required for $\mathsf{CtCtMult}$ and $\mathsf{CtRotate}$ operations.
  • Reducing the cost of BFV operations. Number Theoretic Transformation (NTT) speeds up polynomial multiplications in BFV operations. We implemented NTT using Harvey’s butterfly [38] and optimized it using lazy modular reduction. In every butterfly operation, we track multiplications and only perform modulus when the result overflows the machine word size. This optimization improves forward and backward NTT by $20-25\%$. In BFV multiplication, we perform tensor product and addition in large basis, then scale it back to modulus $Q$. When performing many BFV multiplications, we skip the scaling step after each multiplication and perform it only once at the end [39]. This results in $10-20\%$ improvement in MulPIR used for fetching metadata.
  • Reducing the size of response ciphertexts. At the end of the Wally protocol, the server sends response ciphertexts back to the client. The client just decrypts these ciphertexts and will not use them for further computation. Due to this, we can use the standard modulus switching technique to reduce the size of these ciphertexts by scaling down to the RNS limb $q_{l}$. In BFV after mod switch, the message is encoded into higher $\log_{2}{t}$ bits. Therefore, we can drop the least significant $\log_{2}{q_{l}/t}$ bits of the response ciphertexts and still correctly decrypt the message. However, doing so does add some noise in the resulting ciphertext. We carefully monitor this noise and drop the bits that still allow correct decryption. In [40], the authors suggested that we can drop $l_{0}$ and $l_{1}$ bits from the first and second ciphertext polynomials as long as $z\sqrt{2\frac{n}{9}}2^{l_{1}}+2^{l_{0}}<2^{\log_{2}{(q_{l}/t)}}$, where $z$ represents a score of uniform distribution. The decryption error is bounded with z score probability. We picked $z=8$ that gives decryption error of $2^{-37.5}$. We set $l_{0}=\log_{2}{(q_{l}/t)}-2$, and calculates $l_{1}\leq\lfloor\log_{2}(q/t)-2-\log_{2}(z\sqrt{2n/9})\rfloor$. If $l_{1}\leq 1$ we update $l_{0}=l-1$ and $l_{1}=0$.
    Note that we assume that the original error in the ciphertext is smaller than the error introduced by dropping bits.
  • Reducing the request size. As mentioned in Section II, evaluation keys are essential for SHE operations. In more detail, during the search, the server utilizes evaluation keys to rotate and merge ciphertexts after multiplication. To minimize key size, the client transmits only two keys in two rotations: one for the giant step and another for the baby step (the server repeatedly uses the same keys for all other rotations). Furthermore, we introduce an RNS-based plaintext decomposition method that maintains a constant key size while achieving high precision. Specifically, we select the modulus $t^{\prime}$ (which allows high precision without wrap-around; more details in the next section) as a multiple of two smaller NTT-friendly primes, $t_{0}$ and $t_{1}$. The server decomposes each database vector into two RNS components. The client decomposes the query into two RNS components, encrypts each separately using the same secret key $s$ but different plaintext modulus as $t_{1}$. The server performs exact dot-product computations using each query ciphertext, and the server returns both of them to the client. After decryption, the client reconstructs the original plaintext in $Z_{t}$ using the Chinese Remainder Theorem (CRT). Our main observation is that the evaluation keys do not depend on the plaintext modulus used in the ciphertext; therefore, the server will use the same evaluation key for both queries.
    In metadata fetch, to expand a query, the server performs repeated substitution operations on the query ciphertext. Typically, the server requires one substitution key for each required power. However, we propose a method where the client sends keys for only a few substitutions, and the server computes the remaining substitutions using the same keys. We provide detailed explanations in the Appendix.
VII-B Embedding and Clustering
We map each document in the server database into dense embeddings that retain semantic information. For this, we use the sentence embedding model ‘msmarco-distilbert-base-tas-b’ with 66 million parameters and optimized for semantic search [^3] https://huggingface.co/sentence-transformers/msmarco-distilbert-base-tas-b. In particular, we use the implementation provided by the ‘sentence-transformers’ Python library [^4] https://www.sbert.net. The model outputs dense 768-dimensional floating point vectors. We then use PCA to reduce the dimensionality of each vector to 192.
The next step is clustering these vectors. To do so, we first find $K$ cluster centroids by running k-mean clustering. Then we assign each embedding the cluster with the nearest centroid. Finally, all points and their assigned cluster are compiled into an index.
We implemented this pipeline using the open-source library FAISS. FAISS provides fast, GPU-based algorithms for k-means clustering and indexing. For a comprehensive list of optimizations implemented by FAISS, refer to https://github.com/facebookresearch/faiss/wiki#research-foundations-of-faiss. FAISS exposes a flexible clustering API that allows us to experiment with various clustering parameters, such as the number of data points sampled, the number of iterations of the basic k-means algorithm, and whether to run on GPU or CPU. For indexing, we utilize one of the Flat indexes provided by FAISS to perform fast intra-cluster nearest neighbor search.
VII-C Parameters Selection
SHE parameters.
Performance of Wally directly depends on our choice of BFV parameters. We tried multiple parameter sets and selected the one that yielded the optimal performance. For search and metadata fetch, the selected parameters are as follows: polynomial degree $n=4096$, ciphertext modulus of $\log_{2}{Q}=83$ bits, with CRT limbs of $27,28$  and $28$ bits. We sample error from a centered binomial distribution with standard deviation $3.2$ and secret key from a ternary distribution modulus $Q$. This achieves 128 bits of quantum security.
For search, we set plaintext modulus of $\log_{2}{t}=16$ bits. Recall that the embedding model outputs floating-point vectors of dimension $d$, however BFV operates over integer vectors. We convert each float into a fixed-point integer by multiplying it by the appropriate scaling factor $p$ and then rounding down to the nearest integer. We select $p$ to ensure the dot products do not wrap around modulo $t$. We say that precision is equal to $\log_{2}(p)$ bits. To avoid wrap-around, it is important that $p<\sqrt{(t-1)/2}-\sqrt{d}/2$, which implies that $\log_{2}(p)<1/2\log_{2}t\approx 7$. Therefore, we obtain at least 7 bits of precision. To increase the precision, we use the plaintext CRT decomposition. For this, we pick two plaintext CRT limbs of plaintext modulus of $\log_{2}{t_{0}}=16\text{ and }\log_{2}{t_{1}}=17$ bits. Total plaintext bits are 31 bits, which gives us a precision of 15 bits without increasing evaluation key size.
As mentioned above, we reduce the size of the response ciphertext by first mod switching down to the last CRT limb of 27 bits and then further dropping the least significant bits. To avoid decryption error due to dropping bits, we set $l_{0}=9$ and $l_{1}=0$. This drops the size of the single response ciphertext from 55 KB to 23.5 KB.
For MulPir to handle larger noise growth, we use a plaintext modulus of $\log_{2}{t}=5$ bits. Having a smaller plaintext modulus allows us to set $l_{0}=19$ and $l_{1}=13$, which reduces each response ciphertext to 12 KB.
Epoch and number of users.
As described in Section III, Wally operates in discrete epochs. In each epoch, a total of $U$ users submit queries. However, malicious users can undermine Wally’s privacy guarantees by refraining from submitting fake queries. For our privacy analysis, we assume that half of the total clients in any epoch behave maliciously. We compute the privacy loss by excluding fake queries from these malicious clients. In other words we consider worst-case behavior of all the malicious clients.
Longer epochs lead to a larger $U$, which in turn reduces the number of fake queries required per user. However, extending the epoch duration may negatively impact service quality, since some queries might only be processed at the epoch’s end. For our evaluation, we use a epoch size of one and ten minutes and assume that at least 100,000 and 500,$0$ users submit queries during each epoch, of which half are honest. Given that search engines like Google and Bing process millions of queries per minute and serve billions of users, 500k is a conservative estimate in the context of modern search engines [41].
We also assume that each user will participate in around 400 epochs, one or two epochs each day of a year.
Differential privacy parameters.
For a fixed failure probability $\delta$, a smaller $\epsilon$ offers stronger privacy but compromises performance since each client must execute more fake queries. We set $\delta=2^{-30}$, ensuring that DP will fail only once in a billion queries, and $\epsilon=1/800$, providing a reasonable trade-off between privacy and performance. In each epoch, we assume that each participating client makes $\Delta$ queries. These queries represent the number of nearest clusters fetched for each client’s query. As mentioned in Section VI, this guarantees $(1/400,2^{-26})$ differential privacy per epoch. Consequently, the total differential privacy for each client across 400 epochs is $(1,2^{-26})$.
For each cluster, a client samples fake queries from $\mathcal{NB}(r/U,p)$. Which means in expectation the number of fake queries is
$\mathbb{E}[\text{Fake queries per client}]=\frac{rpK}{(1-p)U}$ (1)
$=\frac{3(1+\log_{2}(1/2^{-30}))e^{-0.0005/\Delta}K}{1-e^{-0.0005/\Delta}U}$ (2)
where $K$ is the total number of clusters and $U$ is the total number of participating honest clients. A little bit work shows that the above equation is always $\leq 102,700\Delta K/U$. As mentioned above considering honest users $U=250,000$, a modest estimate given modern search engines’ scale [41]. Therefore number of fake queries is upper bounded by $\leq 0.205\Delta K$
Search parameters and metric
The product of clusters $K$ and the number of clusters probed $\Delta$ determine the performance and accuracy of Wally. A smaller $K$ and a smaller $\Delta$ both enhance performance by reducing the overhead of fake queries. Further notice that very small $K$ will result in large response size. Similarly, reducing $\Delta$ may lead to a substantial drop in accuracy due to the increased probability of missing the true nearest neighbor.
Our objective is to minimize the fake queries overhead, which is proportional to the product $K\Delta$, while balancing search quality and response size. However, to achieve this using exhaustive search over a candidate set of parameters would require considerable time.
We experimented with different parameters and selected those that provide best accuracy while keeping $K$ moderately small and minimizing $\Delta$ as much as possible. For each probe, the client generates a new query with a large evaluation key. Consequently, large $\Delta$ will result in large total request size.
Dataset
We use the MSMARCO document ranking dataset  [42]. The dataset contains approximately 3.2 million passages, where each passage comprises a document identifier, document URL, document title, and document body.
VIII Evaluation
In this section, we demonstrate Wally’s concrete performance and compare it with baselines.
@@ VIII Evaluation
Metrics
To compare the performance, we use following metrics
  • Queries per second (QPS): this metric reflects how many queries the server can process in a second for fixed infrastructure. We assume that the server has at least 10,000 cores. We measure the latency for a single query on a single core and divide the total number of cores by that number.
  • Bandwidth: We calculate the size of the client query and the server response for each query.
  • Accuracy: To evaluate the accuracy of our search results, we used the Mean Reciprocal Rank (MRR@100) metric, a standard measure for assessing the quality of search results. Given a query $q$ and a ground truth index $i$ such that $DB[i]$ is the most relevant entry, we merge all scores (along with entry indices) obtained from probing $\Delta$ clusters, sort them, and retain the first 100 scores. If the list contains $i$ at the j-th rank (where $j\leq 100$), the reciprocal rank score is calculated as $1/j$. In the absence of $i$ in the list, MRR@100 is set to 0.
Baselines.
We compared Wally with following baselines. Both baselines offer full obliviousness but have high performance overhead:
  • Tiptoe (simulated) [14], which is a state of the art private search system based on linear homomorphic encryption (LWE) . Similar to Wally, Tiptoe uses a clustering approach to find the nearest candidate. However for each query they only probe a single cluster. Unfortunately, due to resource constraints we are unable to run the complete Tiptoe system. We therefore simulate it by using following estimates
    • Latency: For latency we estimate the minimum time the server will take to process a single query. We exclude network cost, which gives advantage to Tiptoe as their network bandwidth is quite high. For each query, the computation is dominated by dot-products between vectors. The server first generates a token by multiplying RLWE encryption of secret key with precomputed LWE ciphertext. For this $16\sqrt{N}$ dot products are performed between vectors of dimension $n=2048$. Note that even though this computation can be performed in offline phase, the server still has to perform it for each query. Then to privately fetch and rank candidates from a particular cluster, the server performs $N$ dot-products between vectors of dimension $d$. For each vector we assume that each element is of 32 bits.
    • Accuracy: To simulate their accuracy results we divide the database into $\sqrt{N}$ clusters, probe a nearest cluster and then calculate MRR@100 of the resulting candidates.
    • Communication: To estimate communication we calculate minimum number of 32 bit elements the client and server have to transfer per query. To generate a token, the client sends at least $n=2048$ RLWE ciphertexts or $n^{2}$ elements. For ranking, the client sends a query consisting of $d\sqrt{N}$ elements. For tokens, the server response consists of $16$ RLWE ciphertexts (each consisting of at least 2048 elements) and for ranking the response consists of $\sqrt{N}$ elements.
  • Pacmann, [43] which is a concurrent work. Like Wally, this scheme also employs sublinear server computation, but it offers full obliviousness. Pacmann’s approach differs from Wallyin two key aspects. Firstly, to locate the nearest neighbor, Pacmann utilize a graph-based search algorithm. Secondly, Pacmann employs sublinear Offline/Online PIR to retrieve graph nodes and traverse the graph.
    For each query, the scheme necessitates multiple round trips and per-query maintenance. We compare our results with the ones provided in the Pacmann paper [43], which is sufficient since we conducted our own experiments in the matching environment.
For Wally, we calculate the QPS and bandwidth by summing the cost of $\Delta$ real queries and fake queries.
Comparison with Tiptoe.
The results are presented in Table II. Wally significantly outperforms Tiptoe across all metrics. Notably, Wally’s QPS is 11~27 times better, this is because for each query the server has to perform at least one expensive operation per entry in the database. Similarly, the communication per query is 10~34 times smaller than Tiptoe, for varying numbers of clusters probed. Tiptoe can only probe a single cluster per query, and to probe more clusters, a new query must be initiated. For $\Delta=1$, a single probe case, Wally’s MRR100 score is only slightly better than Tiptoe’s score of 0.11. However, as we increase $\Delta=5$, Wally’s MRR100 score improves to 0.18, which is considerably better than Tiptoe’s score.
Comparison with Pacmann.
As given in the table, Pacmann’s scheme has better QPS and MRR100 than Wally. Specifically, the better QPS is because Pacmann is based on offline/online sublinear PIR and their MRR100 is better due to use of graph based search data structure which is superior than clustering technique. However, their scheme requires large client storage and communication per query. Specifically, for 3.5 million database, the client is required to store a hint of 614 MB and per query communication is 61.6 MB, which is prohibitively high, specifically for cellular networks. Another downside is that hint-based protocol can not directly handle database updates. Finally, to gain better MRR100 the scheme requires multiple round trips which adds to their network delay. We did not consider the cost of network delay in their QPS calculation. Specifically, due to their use of offline/online PIR, even though QPS of Pacmann is 1.35.6×1.3$\sim$ 5.6$\times$1.3 ∼ 5.6 × better than Wally but the communication is 23123×23$\sim$ 123$\times$23 ∼ 123 × as large for varying $\Delta$.
Client Storage (MB) Communication (MB) Queries Per Second (QPS) MRR100
Tiptoe [14] 0.61 17.4 909 0.11
Pacmann [43] 614 61.6 34,482 0.26
Wally ($K=256$, $\Delta=1$) 0.04 0.56 25,974 0.12
Wally ($K=256$, $\Delta=3$) 0.04 1.7 9,881 0.16
Wally ($K=256$, $\Delta=5$) 0.04 2.6 6,667 0.18
TABLE II: Comparison of Wally  Tiptoe and Pacmann for MSMARCO [42] database. Communication and QPS for Wally includes overhead due to fake queries. Worst value in each metric is marked with a red color.
Benchmarking Wally.
In Table III, we evaluate the performance of Wally across various database sizes ranging from one million entries to hundred million entries and cluster sizes ranging from 128 to 1,024. For all these experiments, we assume that the client probes the nearest cluster $\Delta=1$. Additionally, we consider epoch sizes of one and ten minutes, with 100,000 and 500,000 thousand honest users, respectively.
Firstly, we notice that increasing the epoch size (i.e., the number of honest users) leads to improved overall performance across all database sizes. Specifically, the QPS increases by 2-3 times, while the request and response sizes improve by 1.5-2.4 times. This demonstrates that Wally is an ideal choice for applications where moderate to extended delays are acceptable.
Secondly, we observe that the expected number of fake queries is nearly proportional to the product of the number of honest users ($U$) and inverse of the number of clusters ($K$). For each database, as the epoch size increases from one minute to ten minutes, we increase the number of users by 5×5$\times$5 × (which reduces fake queries) but also increase the number of clusters by 2×2$\times$2 × (which increases fake queries). Consequently, we see that fake queries almost decrease by a factor of $5/2=2.5$, indicating a linear dependency. Alternatively, we can consider that reducing the number of clusters $K$ for a fixed database will lead to a decrease in fake queries.
Thirdly, we observe that as the database size varies, the request and response sizes linearly increase. However, the QPS decreases. This is because for larger databases, each cluster size becomes substantial, resulting in increased server operations per query.
Entries in database N 1 Million 16 Million 100 Million
Epoch size 1 Minute 10 minutes 1 Minute 10 minutes 1 Minute 10 minutes
Number of users $U$ 100,000 500,000 100,000 500,000 100,000 500,000
Number of clusters (K) 128 256 256 512 512 1,024
Expected fake queries 1.7 0.3 3.3 1.7 6.7 2.6
Request size (MB) 0.76 0.36 1.18 0.76 2.17 1.01
Response size (MB) 0.17 0.04 2.3 1.5 13.9 3.03
Queries per second (in thousand) 222 666 14 32 4.7 13.9
TABLE III: Performance of Wally across varying database sizes. Each client will probe a single cluster $\Delta=1$. Each setting provides $(\epsilon=1,\delta=2^{-26})$-DP guarantee.
Number of clusters (K) 64 128 256 512
Expected fake queries 0.3 0.6 1.2 2.4
Request size (MB) 64 128 256 512
Response size (MB) 64 128 256 512
QPS (in thousand) 64 128 256 512
TABLE IV: Performance of Wally for different values of $K$. Each epoch is of ten minutes and database has one million entries. Number clusters probed $\Delta=3$. Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
Clusters probed ($\Delta$) 1 3 5 10
Expected fake queries 0.3 0.6 0.8 1.3
Request size (MB) 0.36 0.45 0.59 0.65
Response size (MB) 0.08 0.10 0.11 0.14
QPS (in thousand) 370 312 277 217
TABLE V: Performance of Wally for different values of $\Delta$. Each epoch is of ten minutes and database has one million entries. Number of clusters $K=125$. Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
IX Related Work
Early private search systems.
Early private search systems heuristically add dummy queries or obfuscate queries by changing the query client-side, but they ultimately send queries in the clear to the server [44, 45, 46, 47, 6, 48, 49, 50, 51]. These systems are broken by a semi-honest server, i.e., the server alone can de-anonymize queries [52, 53, 54, 55] or they have strict distribution requirements on queries, e.g., i.i.d. queries [47]. This contrasts with Wally  which has a provable differential privacy guarantee: the server’s view between the histograms over clusters with and without a client’s queries is differentially private. Further, the client’s queries in Wally are encrypted under homomorphic encryption to prevent the trivial attack shown in [26, Section 3].
Homomorphic encryption.
Homomorphic encryption alone is a popular technique which has been used in private search systems [14, 56, 57, 58]. A major downside of these systems is that they require encrypted computation over the entire database, whereas Wally computes on a few clusters per query. Wally achieves this by utilizing differential privacy to hide the requested cluster index from the server, which makes Wally practically deployable in settings with many honest clients. On the other hand, these previous systems based only on homomorphic encryption have the strongest possible privacy notion: fully obliviousness from the server’s point of view assuming standard cryptographic hardness assumptions. An in-depth comparison with Tiptoe [14] in Section I. A related system is presented in [6] which sends text queries in the clear but uses additively homomorphic encryption [59] to compute an encrypted similarity score. The queries are noised, or “embellished”, beforehand by replacing each word in the query with a set of unrelated words. Embellishing heuristics provide no provable security guarantee, unlike the differential privacy guarantee in Wally.
Secure multiparty computation.
Secure multi-party computation (MPC) is a common technique for private search systems. These systems incur large query sizes or high latency due to many communication rounds. SANNS [60] uses garbled circuits and homomorphic encryption for private nearest neighbors search in the semi-honest client-server model. As a result, they have complete cryptographic obliviousness for each query and the server’s database entries are private besides the query’s response, but each query sends gigabytes of data for databases of 1M entries or more due to their use of garbled circuits [60, Table 1]. Preco [61] is a two-server maliciously-secure MPC private nearest neighbor search system which requires an linearly-sized computation by each server ($N$), a $\sqrt{N}$ computation by the client, and $O(N)$ communication for each query [61, Table 1]. They achieve fully cryptographic obliviousness via distributed point functions [62] and probabilistic batch codes [63]. Concretely, they achieve 10MB of communication and 10s latency for a database with 10M entries.
Differential privacy.